# Mutation-acyclic quivers are totally proper

### v1

@@ Abstract.
Totally proper quivers, introduced by S. Fomin and the author [11], have many useful properties including powerful mutation invariants. We show that every mutation-acyclic quiver (i.e., a quiver that is mutation equivalent to an acyclic one) is totally proper. This yields new necessary conditions for a quiver to be mutation-acyclic.
@@ 1. Introduction
Quivers are finite directed graphs without oriented cycles of length $1$ or $2$. Mutations are operations that transform a quiver, based on a choice of a vertex. These notions are foundational in the study of cluster algebras [15]. A mutation invariant is a characteristic of a quiver that is preserved under mutations. Mutation invariants are helpful for deciding whether two quivers are mutation equivalent or not, i.e., whether there is a sequence of mutations that transforms one quiver into the other. See [5, 9, 14, 20, 22] for examples of known mutation invariants.
@@
In this paper, we prove that for one important class of quivers, the properness requirement can be lifted, so that mutations at all vertices are allowed and all invariants developed in [11] become true mutation invariants. Specifically, we show that in any mutation-acyclic quiver $Q$ (i.e., a quiver that is mutation equivalent to an acyclic quiver), every vertex $v$ is proper, for a particular canonical cyclic ordering $\sigma$ on $Q$. After a mutation at $v$, we obtain a new COQ $(Q^{\prime},\sigma^{\prime})$, with canonical cyclic ordering $\sigma^{\prime}$, that again has the same property: all the vertices are proper, so we can mutate at any one of them, and the process continues.
@@
We have shown in [11, Theorem 11.3, Remark 11.11] that a quiver $Q$ can be upgraded to a totally proper COQ $(Q,\sigma)$ in at most one way (up to wiggles). Moreover, the (essentially unique) candidate cyclic ordering $\sigma=\sigma_{Q}$ can be constructed efficiently. We say that a quiver $Q$ is totally proper if the COQ $(Q,\sigma_{Q})$ is totally proper.
As shown in [11, Corollary 11.2], totally proper quivers have a powerful mutation invariant, which we recall next. This invariant is constructed as follows.
@@
Step 2. “Tear” $\sigma$ into a linear ordering <<\,<.
Step 3. Construct the skew-symmetric exchange matrix $B=B_{Q}$, with rows and columns ordered acording to $<$.
@@
We have shown in [11] that wiggles, cyclic shifts of the linear ordering, and (proper) mutations of COQs preserve the integral congruence class of the unipotent companion. While the resulting invariant of proper mutations is very powerful, it is not easy to use in practice, since the problem of deciding whether two upper-triangular matrices are congruent over ${\operatorname{GL}}_{n}({\mathbb{Z}})$ seems to be rather difficult. In [11], we bypassed this difficulty as follows. It is well known (and easy to see) that the integral congruence class of a matrix $U\in{\operatorname{GL}}_{n}({\mathbb{Z}})$ uniquely determines the conjugacy class of its cosquare $U^{-T}U$. It follows that whenever two COQs are related by proper mutations, the cosquares of their respective unipotent companions must be conjugate in ${\operatorname{GL}}_{n}({\mathbb{Z}})$. This conjugacy condition can be verified efficiently, though the algorithm is quite nontrivial [3, 7].
@@
The main result of this paper is the following theorem that settles [11, Conjecture 12.8].
@@
Any mutation-acyclic quiver is totally proper.
@@
It suffices to show that any acyclic quiver is totally proper. Furthermore, as noted in [11, Observation 10.7], any acyclic quiver has a proper cyclic ordering obtained by “closing up” any linear ordering compatible with the orientations of the quivers’ arrows. (All such cyclic orderings are related by wiggles.) Our proof of Theorem 1.1 shows that these cyclic orderings are totally proper.
In the course of proving Theorem 1.1, we establish a few results that may be of independent interest. In particular, Theorem 4.4 gives a simple combinatorial constraint on the orientations of arrows in mutation-acyclic quivers. Both Theorem 4.4 and the ensuing Lemma 5.5 assert that certain subquivers of a mutation-acyclic quiver must be acyclic. Should someone wish to compute with COQs and unipotent companions, Lemmas 4.1 and 4.2 generalize and remove conditions from [11, Lemma 7.2 and Remark 7.3]. This can speed up and simplify the computations.
@@
Section 2 establishes notation and reviews some material that does not involve COQs. Much of this section is devoted to restating the results of A. Seven [21], who constructed a distinguished symmetric matrix $A$ associated with an arbitrary mutation-acyclic quiver. Section 3 is a condensed summary of the necessary results and notation from [11]. It contains new examples, and Proposition 3.21 gives an alternative description of the Markov invariant. In Section 4 we establish a handful of technical lemmas. Lemmas 4.1 and 4.2 make it easier to compute the unipotent companion of a COQ after a proper mutation. Theorem 4.4 generalizes [11, Corollary 10.12] using the results of A. Seven [21]. Section 5 contains the proof of Theorem 1.1. We begin by defining $U=(A-B)/2$, where $A$ comes from Seven’s construction and $B$ is the usual exchange matrix. The fact that the integral congruence class of $U$ is a mutation invariant follows quickly from Proposition 2.30. The bulk of the effort is then spent showing that $U$ is indeed a unipotent companion. In Section 6, we give some applications and corollaries of Theorem 1.1. This includes a description of invariants for some small quivers, inequalities satisfied by the coefficients of the Alexander polynomial for mutation-acyclic quivers (Corollary 6.6), and several examples of quivers that are shown to be not mutation-acyclic.
@@ Acknowledgements
I would like to thank my advisor, Sergey Fomin, for his guidance, comments, and suggestions; Grayson Moore, for reading and commenting on an early draft; and Danielle Ensign for her software assistance. I am also grateful to Roger Casals, Tucker Ervin, and Hugh Thomas for stimulating discussions.
@@ 2. Quivers, mutation, and quasi-Cartan companions
Figure 2. Two quivers on $3$-vertices. The quiver on the left is not mutation-acyclic, while the quiver on the right is mutation-acyclic (by mutating at $v_{2}$ and then $v_{1}$).
@@
For a quiver $Q$, the $B$-matrix $B_{Q}=(b_{uv})$ (also known as the exchange matrix) is the skew-symmetric adjacency matrix of $Q$. By convention we have $b_{uv}>0$ when there are $b_{uv}$ arrows $u\rightarrow v$, and $b_{uv}<0$ when there are $-b_{uv}$ arrows $u\leftarrow v$.
@@
Fix a quiver $Q_{0}$ with $n$ mutable linearly ordered vertices $v_{1}<\cdots<v_{n}$ (and no frozen vertices). We index the mutation class $[\widehat{Q}_{0}]$ by simultaneously assigning a quiver $\tilde{Q}_{t}\in[\widehat{Q}_{0}]$ to each vertex $t$ in $\mathbb{T}_{n}$, so that $\tilde{Q}_{t_{0}}=\widehat{Q}_{0}$ and $\tilde{Q}_{t}=\mu_{v_{i}}(\tilde{Q}_{t^{\prime}})$ whenever $t\stackrel{{\scriptstyle i}}{{\text{---}}}t^{\prime}$. We call $Q_{0}$ the initial quiver. For each vertex $t\in\mathbb{T}_{n}$, let $Q_{t}$ be the mutable part of $\tilde{Q}_{t}$. Let $B_{t}=(b_{v_{i}v_{j};t})$ be the $n\times n$ exchange matrix of $Q_{t}$. Similarly, let $C_{t}=(c_{v_{i}^{\prime}v_{j};t})$ be $n\times n$ bipartite adjacency matrix between frozen vertices and mutable vertices. Thus
@@
The map $t\mapsto(C_{t},B_{t})$ is called a (principal, tropical) $Y$-seed pattern in [21], as it is essentially equivalent to an assignment of tropical coefficient variables (usually denoted $\mathbb{y}$) and $B$-matrix to each vertex $t\in\mathbb{T}_{n}$, cf. [14, Proposition 3.6.5].
@@
  • for each red mutable vertex $v_{j}$, add $-\sum_{i}c_{v_{i}^{\prime}v_{j};t}$ arrows $\mathring{v}\leftarrow v_{j}$.
@@
More generally, $C$-matrices can be used to compute the number of arrows between frozen and mutable vertices for “triangular extensions” other than the principal framing, such as those that involve adding a frozen vertex $\mathring{v}$ and any number of arrows $\mathring{v}\rightarrow v$, cf. [4, Theorem 3.2].
@@
For a quiver $Q$ we define the underlying unoriented simple graph $K_{Q}$. The graph $K_{Q}$ has the same vertex set as $Q$, and an edge $u-v$ whenever there are any arrows $u\rightarrow v$ or $v\rightarrow u$ in $Q$. There are no parallel edges.
@@
A quasi-Cartan companion $A=(a_{vw})$ of a quiver $Q$ is admissible if for every chordless cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$ the number #$\#\{i|a_{w_{i}w_{i+1}}>0\text{ and }i<\ell\}$ is odd when $\mathcal{O}$ is oriented, and is even otherwise.
@@
Theorem 2.26 ([21, Theorems 1.3 and 1.4]).
Let $Q_{0}$ be acyclic, and fix $t\in\mathbb{T}_{n}$. Let $A_{t}=(a_{uv;t})$.
@@
To see how [21, Theorem 1.3] implies the first statement of Theorem 2.26, assume that $b_{ji}<0$ and check the four combinations of $\operatorname{sgn}(\mathbb{c}_{i})$ and $\operatorname{sgn}(\mathbb{c}_{j})$.
@@
For quivers mutation equivalent to acyclic quivers with sufficiently many arrows, a number of properties of the matrices $B_{t}$ and $C_{t}$ matrices have been recently established by T. Ervin, see [8, Proposition 5.2, Lemma 5.4]. These results may potentially be used to obtain enhancements of Theorem 2.26.
Definition 2.29 ([14, Theorem 2.8.3]).
We define a pair of square matrices $M_{Q}(v,\pm 1)$, associated to mutating a quiver $Q$ at a given vertex $v$. Let ${B_{Q}=(b_{pq})}$ be the exchange matrix of $Q$ and choose $\varepsilon=\pm 1$. Let $J$ be the diagonal matrix with $J_{vv}=-1$ and $J_{qq}=1$ for $w\neq v$. Let $E=(e_{pq})$ be the matrix with $v$th column $e_{qv}=\max(0,-\varepsilon b_{qv})$ and 00 otherwise. Then $M_{Q}(v,\varepsilon)=J+E$.
@@
Proposition 2.30 ( [14, Theorem 2.8.3], cf. [21, Lemma 3.1]).
@@ 3. Cyclically ordered quivers
Figure 5. The $4!/4=6$ distinct COQs on the $4$-vertex acyclic quiver with vertices and arrows $v_{1}\rightarrow v_{2}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{4}$, $v_{1}\rightarrow v_{3}\rightarrow v_{4}$, and $v_{1}\stackrel{{\scriptstyle 3}}{{\rightarrow}}v_{4}$.
@@
The unipotent companion of a linearly ordered quiver $Q$ is the unipotent upper triangular matrix $U$ such that $U^{T}-U=B_{Q}$, where the rows and columns of $B_{Q}$ are written in the linear ordering. The matrix $U$ can be constructed by negating $B_{Q}$, then setting the lower triangular part to 00 and the diagonal entries to $1$. For a COQ $Q$, the unipotent companion of any associated linearly ordered quiver is a unipotent companion of $Q$.
@@
Definition 3.8 (Cf. [11, Definition 2.10]).
Fix a COQ $(Q,\sigma)$. Let $\mathcal{O}=(w_{0}-w_{1}-\cdots-w_{k}=w_{0})$ be a cycle in $K_{Q}$ (Definition 2.20). Fix a linear order $<$ associated to $\sigma$. The winding number ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ is the (signed) number of times we “wrap around” the linear order:
@@
We omit the (straightforward) proofs that ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ does not depend on the choice of linear order associated to $\sigma$, and that this definition agrees with [11, Definition 2.10].
@@
Let $Q$ be the underlying quiver of the COQs shown in Figure 6, and consider the (chordless) cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{1})$ in $K_{Q}$. The winding number of this undirected cycle is 00 in the three COQs in the top row and $1$ in the three COQs in the bottom row.
@@
On the left, a COQ with quiver whose arrows are v1 to v2, v2 to v4, v1 to v3, v3 to v4, and whose cyclic ordering is (v1, v2, v3, v4). On the right, the same vertices drawn on a dashed circle in the same positions as the COQ, but the arrows have been drawn along the dashed circle. The arrow from v1 to v2 now lies on the circular arc clockwise from the location of v1 (135 degrees on the circle) to the location of v2 (45 degrees). The arrow from v2 to v3 lies on the circular arc clockwise from the location of v2 to the location of v4 (-135 degrees). The arrow from v3 to v4 lies on the circular arc clockwise from the location of v3 (-45 degrees) to the location of v4. The arrow from v1 to v3 lies on the circular arc clockwise from the location of v1 to the location of v3. In particular, none of the arrows go through the point at 180 degrees, while they cover each point from 135 degrees to -135 degrees (traveling clockwise) twice.
Figure 7. A COQ $Q$ and the image of the cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{3}-v_{1})$ in $K_{Q}$, with the arrows drawn on the circle. In this case the image is contractible, and so the winding number is 00.
Theorem 3.11 ([11, Theorem 2.12]).
@@
A sequence of three proper mutations between four COQs. Starting on the left, the quiver has vertices v1, v2, v3, v4 (with that cyclic ordering), and an arrow from v2 to v1, v3 to v1, v3 to v4, as well as two arrows from v2 to v4 and 3 arrows from v4 to v1. In particular, v1 is a sink vertex. Thus there are no oriented paths through v1, and so v1 is proper. The first mutation is at v1, which only reverses the arrows touching v1, and leaves the cyclic ordering unchanged. Notice that the only oriented path of length two through v2 starts at v1 and ends at v4. This is a right turn, so v2 is proper. We mutate at v2 to arrive at a new COQ with cyclic ordering (v1, v3, v4, v2) and an arrow from v1 to v3, v3 to v4, v2 to v1, as well as five arrows from v1 to v4 and two arrows from v4 to v2. As v2 had an oriented path through it, we have changed its position in the cyclic ordering. We next wish to mutate at v3. It is proper as the only oriented 2-path goes from v1 to v4, and v1, v3, v4 is a right turn. Proper mutation at v3 results in the COQ with cyclic ordering (v1, v4, v3, v2) and a single arrow from v2 to v1, v3 to v1, v4 to v3, as well as two arrows from v4 to v2 and six arrows from v1 to v4.
@@
Lemma 3.18 ([11, Lemma 7.2 and Remark 7.3]).
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$ and $v_{i}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.

### v2

@@ Abstract.
Totally proper quivers, introduced by S. Fomin and the author [13], have many useful properties including powerful mutation invariants. We show that every mutation-acyclic quiver (i.e., a quiver that is mutation equivalent to an acyclic one) is totally proper. This yields new necessary conditions for a quiver to be mutation-acyclic. In particular, we show that a generalization of the Markov invariant for $3$-vertex quivers applies to all mutation-acyclic quivers. Only finitely many acyclic quivers share the same Markov invariant.
@@ 1. Introduction
Quivers are finite directed graphs without oriented cycles of length $1$ or $2$. Mutations are operations that transform a quiver, based on a choice of a vertex. These notions are foundational in the study of cluster algebras [17]. A mutation invariant is a characteristic of a quiver that is preserved under mutations. Mutation invariants are helpful for deciding whether two quivers are mutation equivalent or not, i.e., whether there is a sequence of mutations that transforms one quiver into the other. One example is the Markov invariant of a $3$-vertex quiver, which essentially characterizes the $3$-vertex quivers which are mutation-acyclic (i.e., those quivers which are mutation equivalent to an acyclic quiver) [2]. See [7, 11, 16, 23, 25] for additional examples of known mutation invariants.
@@
In this paper, we prove that for one important class of quivers, the properness requirement can be lifted, so that mutations at all vertices are allowed and all invariants developed in [13] become true mutation invariants. Specifically, we show that in any mutation-acyclic quiver $Q$, every vertex $v$ is proper for a particular canonical cyclic ordering $\sigma$ on $Q$. After a mutation at $v$, we obtain a new COQ $(Q^{\prime},\sigma^{\prime})$, with canonical cyclic ordering $\sigma^{\prime}$, that again has the same property: all the vertices are proper, so we can mutate at any one of them, and the process continues.
@@
We have shown in [13, Theorem 14.3] that a quiver $Q$ can be upgraded to a totally proper COQ $(Q,\sigma)$ in at most one way (up to wiggles). Moreover, the candidate cyclic ordering $\sigma=\sigma_{Q}$ can be constructed efficiently [13, Algorithm 14.12]. We say that a quiver $Q$ is totally proper if the COQ $(Q,\sigma_{Q})$ is totally proper.
As shown in [13, Corollary 14.2], totally proper quivers have a powerful mutation invariant, which we recall next. This invariant is constructed as follows.
@@
Step 2. “Tear” $\sigma$ into a linear ordering <<\,.
Step 3. Construct the skew-symmetric exchange matrix $B=B_{Q}$, with rows and columns ordered according to <<.
@@
We have shown in [13] that wiggles, cyclic shifts of the linear ordering, and (proper) mutations of COQs preserve the integral congruence class of the unipotent companion. While the resulting invariant of proper mutations is very powerful, it is not easy to use in practice, since the problem of deciding whether two upper-triangular matrices are congruent over ${\operatorname{GL}}_{n}({\mathbb{Z}})$ seems to be difficult. In [13], we bypassed this difficulty as follows. It is well known (and easy to see) that the integral congruence class of a matrix $U\in{\operatorname{GL}}_{n}({\mathbb{Z}})$ uniquely determines the conjugacy class of its cosquare $U^{-T}U$. It follows that whenever two COQs are related by proper mutations, the cosquares of their respective unipotent companions must be conjugate in ${\operatorname{GL}}_{n}({\mathbb{Z}})$. This conjugacy condition can be verified efficiently, though the algorithm is nontrivial [5, 9].
@@
The main result of this paper is the following theorem [13, Conjecture 15.9].
@@
Every mutation-acyclic quiver is totally proper.
@@
It suffices to show that every acyclic quiver is totally proper. Furthermore, as noted in [13, Observation 11.6], any acyclic quiver has a proper cyclic ordering obtained by “closing up” any linear ordering compatible with the orientations of the quivers’ arrows. (All such cyclic orderings are related by wiggles.) Our proof of Theorem 1.1 shows that these cyclic orderings are totally proper.
In the course of proving Theorem 1.1, we establish a few results that may be of independent interest. In particular, Theorem 4.4 gives a simple combinatorial constraint on the orientations of arrows in mutation-acyclic quivers. Both Theorem 4.4 and the ensuing Lemma 5.5 assert that certain subquivers of a mutation-acyclic quiver must be acyclic. Should someone wish to compute with COQs and unipotent companions, Lemmas 4.1 and 4.2 generalize and remove conditions from [13, Lemma 7.2 and Remark 7.3]. This can speed up and simplify the computations.
@@
One consequence of Theorem 1.1 is the existence of a Markov invariant for all mutation acyclic quivers. We establish some basic inequalities for its coefficients and give a simple combinatorial formulation in the acyclic case (Proposition 3.21), including that the Markov invariant is positive (respectively, at least n1n-1) for any (respectively, any connected) mutation-acyclic quiver (Corollary 6.6). As in the $3$-vertex case, there are only finitely many acyclic quivers with a given Markov invariant. Thus there are only finitely many acyclic quivers with a given Alexander polynomial (Corollary 6.7).
Amanda Schwartz [21] studied the Alexander (and HOMFLY) polynomials of the link associated to certain plabic graphs, which in turn correspond to tree quivers (those whose underlying undirected graph is a tree). Our constructions agree on the Alexander polynomial. She gave a combinatorial description of all of the coefficients of the Alexander polynomial in this case [21, Theorem 4.2].
A quasi-Cartan companion is a symmetric matrix associated to a quiver, in analogy to the Cartan companion of a dynkin diagram. The existence of a quasi-Cartan companion with various properties has been used to characterize quivers of finite type [4, 22]. Our proof of Theorem 1.1 uses the results of A. Seven [24] to identify the correct cyclic ordering to associate with a given mutation-acyclic quiver, and to guarantee that certain subquivers of a mutation-acyclic quiver are acyclic (Theorem 4.4). Some of our technical lemmas also extend to quasi-Cartan companions, see Corollary 6.18.
@@
Section 2 establishes notation and reviews some material that does not involve COQs. Much of this section is devoted to restating the results of A. Seven [24]. Section 3 is a condensed summary of the necessary results and notation from [13]. It contains new examples, and Proposition 3.21 gives an alternative description of the Markov invariant for acyclic quivers.
In Section 4 we establish a handful of technical lemmas. Lemmas 4.1 and 4.2 make it easier to compute the unipotent companion of a COQ after a proper mutation. Theorem 4.4 generalizes [13, Corollary 13.3] using the results of A. Seven [24]. Section 5 contains the proof of Theorem 1.1. We begin by defining U=(AB)/2U=(A-B)/2, where $A$ comes from Seven’s construction and $B$ is the usual exchange matrix. The fact that the integral congruence class of $U$ is a mutation invariant follows quickly from Proposition 2.30. The bulk of the effort is then spent showing that $U$ is indeed a unipotent companion.
In Section 6, we give some applications and corollaries of Theorem 1.1. This includes a description of invariants for some particular families of quivers, inequalities satisfied by the coefficients of the Alexander polynomial for mutation-acyclic quivers (Corollaries 6.66.7, and 6.8), and several examples of quivers that are shown to be not mutation-acyclic. This section also contains Examples 6.16 and 6.17, which respectively show that not all Alexander polynomials of acyclic quivers are unimodal, but that many are.
@@ Acknowledgements
I would like to thank my advisor, Sergey Fomin, for his guidance, comments, and suggestions; Grayson Moore, for reading and commenting on an early draft; and Danielle Ensign for her software assistance. I am also grateful to Roger Casals, Tucker Ervin, Hugh Thomas, and Sara Billey for stimulating discussions.
@@ 2. Quivers, mutation, and quasi-Cartan companions
Figure 2. Two quivers on 33-vertices. The quiver on the left is not mutation-acyclic, while the quiver on the right is mutation-acyclic (by mutating at $v_{2}$ and then $v_{1}$).
@@
For a quiver $Q$, the $B$-matrix $B_{Q}=(b_{uv})$ (also known as the exchange matrix) is the skew-symmetric adjacency matrix of $Q$. By convention we have $b_{uv}>0$ when there are $b_{uv}$ arrows $u\rightarrow v$, and $b_{uv}<0$ when there are $b_{uv}$ arrows $u\leftarrow v$.
@@
Fix a quiver $Q_{0}$ with $n$ mutable linearly ordered vertices $v_{1}<\cdots<v_{n}$ (and no frozen vertices). We index the mutation class $\widehat{Q}_{0}]$ by simultaneously assigning a quiver $\tilde{Q}_{t}\in[\widehat{Q}_{0}]$ to each vertex $t$ in $\mathbb{T}_{n}$, so that $\tilde{Q}_{t_{0}}=\widehat{Q}_{0}$ and $\tilde{Q}_{t}=\mu_{v_{i}}(\tilde{Q}_{t^{\prime}})$ whenever $t\stackrel{{\scriptstyle i}}{{\text{---}}}t^{\prime}$. We call $Q_{0}$ the initial quiver. For each vertex $t\in\mathbb{T}_{n}$, let $Q_{t}$ be the mutable part of $\tilde{Q}_{t}$. Let $B_{t}=(b_{v_{i}v_{j};t})$ be the $n\times n$ exchange matrix of $Q_{t}$. Similarly, let $C_{t}=(c_{v_{i}^{\prime}v_{j};t})$ be $n\times n$ bipartite adjacency matrix between frozen vertices and mutable vertices. Thus
@@
The map $t\mapsto(C_{t},B_{t})$ is called a (principal, tropical) $Y$-seed pattern in [24], as it is essentially equivalent to an assignment of tropical coefficient variables (usually denoted $\mathbb{y}$) and $B$-matrix to each vertex $t\in\mathbb{T}_{n}$, cf. [16, Proposition 3.6.5].
@@
  • for each red mutable vertex $v_{j}$, add $\sum_{i}c_{v_{i}^{\prime}v_{j};t}$ arrows $\mathring{v}\leftarrow v_{j}$.
@@
More generally, $C$-matrices can be used to compute the number of arrows between frozen and mutable vertices for “triangular extensions” other than the principal framing, such as those that involve adding a frozen vertex $\mathring{v}$ and any number of arrows $\mathring{v}\rightarrow v$, cf. [6, Theorem 3.2].
@@
For a quiver $Q$ we define the underlying unoriented simple graph $K_{Q}$. The graph $K_{Q}$ has the same vertex set as $Q$, and an edge uvu-v whenever there are any arrows $u\rightarrow v$ or $v\rightarrow u$ in $Q$. There are no parallel edges.
@@
[…]
@@
Theorem 2.26 ([24, Theorems 1.3, 1.4]).
Let $Q_{0}$ be acyclic, and fix a vertex $t\in\mathbb{T}_{n}$. Let $A_{t}=(a_{uv;t})$.
@@
To see how [24, Theorem 1.3] implies the first statement of Theorem 2.26, assume that $b_{ji}<0$ and check the four combinations of $\operatorname{sgn}(\mathbb{c}_{i})$ and $\operatorname{sgn}(\mathbb{c}_{j})$.
@@
For quivers mutation equivalent to acyclic quivers with sufficiently many arrows, a number of properties of the matrices $B_{t}$ and $C_{t}$ matrices have been recently established by T. Ervin, see [10, Proposition 5.2, Lemma 5.4]. These results may potentially be used to obtain enhancements of Theorem 2.26.
Definition 2.29 ([16, Theorem 2.8.3]).
We define a pair of square matrices $M_{Q}(v,\pm 1)$, associated to mutating a quiver $Q$ at a given vertex $v$. Let $B_{Q}=(b_{pq})$} be the exchange matrix of $Q$ and choose $\varepsilon=\pm 1$. Let $J$ be the diagonal matrix with $J_{vv}=-1$ and $J_{qq}=1$ for $w\neq v$. Let $E=(e_{pq})$ be the matrix with $v$th column $e_{qv}=\max(0,-\varepsilon b_{qv})$ and 0 otherwise. Then $M_{Q}(v,\varepsilon)=J+E$.
@@
Proposition 2.30 ( [16, Theorem 2.8.3], cf. [24, Lemma 3.1]).
@@ 3. Cyclically ordered quivers
Figure 5. The $4!/4=6$ distinct COQs on the $4$-vertex acyclic quiver with vertices and arrows $v_{1}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{2}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{4}$, $v_{1}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{3}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{4}$, and $v_{1}\stackrel{{\scriptstyle 3}}{{\rightarrow}}v_{4}$.
@@
The unipotent companion of a linearly ordered quiver $Q$ is the unipotent upper triangular matrix $U$ such that $U^{T}-U=B_{Q}$, where the rows and columns of $B_{Q}$ are written in the linear ordering. The matrix $U$ can be constructed by negating $B_{Q}$, then setting the lower triangular part to 0 and the diagonal entries to $1$. For a COQ $Q$, the unipotent companion of any associated linearly ordered quiver is a unipotent companion of $Q$.
@@
Definition 3.8.
Fix a COQ $(Q,\sigma)$. Let $\mathcal{O}=(w_{0}-w_{1}-\cdots-w_{k}=w_{0})$ be a cycle in $K_{Q}$ (Definition 2.20). Fix a linear order << associated to $\sigma$. The winding number ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ is the (signed) number of times we “wrap around” the linear order:
@@
We omit the (straightforward) proofs that ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ does not depend on the choice of linear order associated to $\sigma$, and that this definition agrees with [13, Definition 2.10].
@@
Let $Q$ be the underlying quiver of the COQs shown in Figure 6, and consider the (chordless) cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{1})$ in $K_{Q}$. The winding number of this undirected cycle is 0 in the three COQs in the top row and $1$ in the three COQs in the bottom row.
@@
On the left, a COQ with quiver whose arrows are v1 to v2, v2 to v4, v1 to v3, v3 to v4, and whose cyclic ordering is (v1, v2, v3, v4).
On the right, the same vertices drawn on a dashed circle in the same positions as the COQ, but the arrows have been drawn along the dashed circle. The arrow from v1 to v2 now lies on the circular arc clockwise from the location of v1 (135 degrees on the circle) to the location of v2 (45 degrees). The arrow from v2 to v3 lies on the circular arc clockwise from the location of v2 to the location of v4 (-135 degrees). The arrow from v3 to v4 lies on the circular arc clockwise from the location of v3 (-45 degrees) to the location of v4. The arrow from v1 to v3 lies on the circular arc clockwise from the location of v1 to the location of v3. In particular, none of the arrows go through the point at 180 degrees, while they cover each point from 135 degrees to -135 degrees (traveling clockwise) twice.
Figure 7. A COQ $Q$ and the image of the cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{3}-v_{1})$ in $K_{Q}$, with the arrows drawn on the circle. In this case the image is contractible, and so the winding number is 0.
Theorem 3.11 ([13, Theorem 2.14]).
@@
A sequence of three proper mutations between four COQs. Starting on the left, the quiver has vertices v1, v2, v3, v4 (with that cyclic ordering), and an arrow from v2 to v1, v3 to v1, v3 to v4, as well as two arrows from v2 to v4 and 3 arrows from v4 to v1. In particular, v1 is a sink vertex. Thus there are no oriented paths through v1, and so v1 is proper.
$\stackrel{{\scriptstyle\mu_{v_{1}}}}{{\longleftrightarrow}}$ The first mutation is at v1, which only reverses the arrows touching v1, and leaves the cyclic ordering unchanged. Notice that the only oriented path of length two through v2 starts at v1 and ends at v4. This is a right turn, so v2 is proper. $\stackrel{{\scriptstyle\mu_{v_{2}}}}{{\longleftrightarrow}}$ We mutate at v2 to arrive at a new COQ with cyclic ordering (v1, v3, v4, v2) and an arrow from v1 to v3, v3 to v4, v2 to v1, as well as five arrows from v1 to v4 and two arrows from v4 to v2. As v2 had an oriented path through it, we have changed its position in the cyclic ordering. $\stackrel{{\scriptstyle\mu_{v_{3}}}}{{\longleftrightarrow}}$ We next wish to mutate at v3. It is proper as the only oriented 2-path goes from v1 to v4, and v1, v3, v4 is a right turn. Proper mutation at v3 results in the COQ with cyclic ordering (v1, v4, v3, v2) and a single arrow from v2 to v1, v3 to v1, v4 to v3, as well as two arrows from v4 to v2 and six arrows from v1 to v4.
@@
Lemma 3.18 ([13, Lemma 7.2 and Remark 7.3]).
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order << such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$ and $v_{i}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.